# CoTAL: Human-in-the-Loop Prompt Engineering, Chain-of-Thought Reasoning, and Active Learning for Generalizable Formative Assessment Scoring

### v2

@@ IV-A Chain-of-Thought Prompting + Active Learning (CoTAL)
CoTAL is structured into three distinct phases, as shown in Figure 3. In Phase I (yellow), teachers and researchers collaborate using ECD to design formative assessments and rubrics that align with curricular goals. The rubrics are developed to evaluate students’ responses to each formative assessment. Phase I is human-driven, with both parties (teachers and researchers) working collaboratively to create the formative assessments and their corresponding grading rubrics. The LLM is not considered in Phase I. This study’s Phase I formative assessments and rubrics are detailed in Section III.
In Phase II (blue), researchers work with the LLM to develop and optimize an initial prompt by (1) sampling a subset of student responses and conducting IRR to establish scoring consensus, (2) employing few-shot ICL and CoT reasoning to align the LLM with human consensus, and (3) iteratively refining the prompt through active learning by generating scores and explanations using a validation set to ensure the LLM’s generations align with human scoring preferences. These steps correspond to the blue diamonds, which represent Response Scoring, Prompt Development, and Active Learning in Figure 3. This procedure was applied to the Science Concepts and Reasoning Task in previous work [12], as well as the three new integrated assessments in science, computing, and engineering discussed in this paper (see Section III).
@@
In Step 1, Response Scoring, human experts independently score a random subset of student responses using the rubrics developed in Phase I and establish their initial degree of agreement based on an IRR metric (e.g., Cohen’s Kappa [36]). When discrepancies arise between the scores assigned by the two reviewers, the underlying reasons for these differences are analyzed and resolved to reach a final consensus. Difficult-to-resolve discrepancies, termed “sticking points,” receive special attention during Prompt Development (Phase II, Step 2) to refine the prompt structure and guide the LLM toward alignment with the human scorers. Sticking points often emerge from “edge cases” where reviewers interpret the same rubric differently. For example, one Engineering Task sticking point was whether a general understanding of design constraints (i.e., the student did not explicitly mention any constraint by name) qualified for 2 points (both reviewers ultimately agreed this was acceptable). This process—randomly sampling the data, calculating Cohen’s $k$, and documenting reasons for disagreement—is repeated until $k\geq 0.70$ to ensure consensus. This is depicted as the “IRR” self-loop in Figure 3.
@@
In Step 2, Prompt Development, the LLM is first given task instructions, followed by the formative assessment question, grading rubric, and additional context to guide the LLM’s scoring. This supplementary information helps the LLM connect student responses to the rubric and comprehend the intended concepts, questions, and scoring criteria. Additionally, various prompt patterns are employed to guide the LLM during inference and help structure its output, including a: (1) persona pattern that assigns the LLM a “persona,” or role, to play while generating its output; (2) context manager pattern that defines the context the LLM should consider while generating its output; (3) template pattern that provides the LLM with a structured output template; and (4) meta language creation pattern that creates a custom language for the LLM to understand (e.g., a textual shorthand notation to help the LLM interpret graphs) [37]. For example, the context manager pattern helped the LLM understand the relationship between the Rules and Debugging Tasks (i.e., the students often referred to conditional statements by “rule” number during the Debugging Task), and the meta language creation pattern allowed us to distill the Debugging Task’s computational model image into LLM-readable text.
@@
After constructing the prompt, labeled few-shot instances, along with CoT reasoning chains, are added to it to align the LLM with the consensus reached by human scorers. Each formative assessment prompt features two types of few-shot examples: (1) ground truth instances, where both scorers agree during Response Scoring; and (2) sticking point instances, where scorers disagree during Response Scoring and the reasons for disagreement carry over to other instances, leading to similar disagreements among the scorers. For sticking point instances, CoT reasoning chains help clarify potential misunderstandings and guide the LLM toward human consensus. All few-shot instances are accompanied by CoT reasoning chains that loosely (i.e., not verbatim) follow this template:
@@
Once the prompt is constructed, Active Learning (Step 3) tests the prompt against a validation set. The LLM’s generations are analyzed by first isolating the incorrectly scored instances. For each subscore that the LLM predicts incorrectly, we identify a “scoring trend” to determine whether the LLM tends to produce false positives (FPs) or false negatives (FNs) and qualitatively discern the reasons behind these inaccuracies. This process mirrors Response Scoring, where recurring LLM errors on the validation set are identified as sticking points. We then select validation set instances exemplifying these sticking points, annotating them with CoT reasoning chains to correct the LLM’s mistakes and adding them to the existing few-shot examples in the prompt (illustrated by the “AL Prompt Refinements” loop in Figure 3).
@@
Phase II of CoTAL shares similarities with explainable AI (XAI) approaches, which focus on explaining a model’s decisions based on its internal mechanisms. However, in CoTAL’s case, the explainability comes from the LLM’s generations, offering insights into how the model evaluates student responses in relation to rubrics. The emphasis is on alignment with grading criteria rather than interpreting the model’s internal logic. This explainability is particularly valued by students, who see it as a critical factor for building trust in AI systems in educational settings (see Section VI). Unlike our prior work, which primarily highlighted our human-in-the-loop prompt engineering contributions, this paper “closes the loop” (see Figure 3) by (1) integrating LLM prompt engineering with ECD principles, (2) investigating our method’s practical utility in classrooms through studies involving both teachers and students, and (3) leveraging student and teacher feedback to inform refinements to our formative assessments, rubrics, and prompts.
@@
In this work, we refine our original approach based on prior findings [12]. Previously, we conducted active learning by inserting several validation instances back into the prompt. However, we found that the LLM tended to overfit when the number of few-shot instances was large, or the CoT reasoning chains became too granular. To mitigate the LLM’s tendency to overfit, we insert only a single instance into the prompt during Active Learning in this study, specifically targeting the most persistent LLM errors. Additionally, we use Prompt Development (Phase II, Step 2) to provide the LLM with a list of “guidelines” [^2] In the actual prompt, we refer to the guidelines as “rules,” but we use the term “guidelines” in the manuscript so as not to be confused with the “rules” in the Rules Task. that the LLM is instructed to adhere to at all times. For example, one Rules Task guideline instructed the LLM that the order in which students listed the three rules should not affect its scoring decisions. Human scorers agree upon these guidelines based on the consensus they reach during Response Scoring. Just as we use the CoT reasoning chains to cite evidence from the student’s response and tie that evidence to the rubric to elicit a score, we similarly use the CoT reasoning chains to cite these guidelines to improve LLM responses.
@@
In Phase III, the prompt refined during Phase II is deployed in a classroom setting to score students’ formative assessment responses and explain the assigned scores. Researchers then sample the LLM’s responses and present them to teachers, who critique the model’s scoring accuracy, explanatory soundness, and clarity. The teachers and researchers identify the LLM’s strengths and weaknesses during these discussions. Based on this feedback, they collaborate to determine how to best address the LLM’s shortcomings without compromising its strengths, agreeing on specific refinements to the formative assessment questions, scoring rubrics, and prompts.
@@
Similarly, students are shown the LLM’s responses to their formative assessment answers and are asked to critique them. While teacher feedback informs methodological improvements related to curricular goals, student feedback highlights user experience and personalized learning. For example, students might emphasize elements such as the LLM’s tone in its responses and the overall effectiveness of the content it provides to enhance their understanding of relevant concepts. This process is illustrated in Figure 3 by the “Curricular and Instructional Refinements” and “Teacher/Student Prompt Refinements” loops. Like Phase I, Phase III depends on human input, involving close collaboration among researchers, students, and teachers to guide curricular and methodological enhancements.
IV-B Experimental Design
@@
We analyzed formative assessment responses from 175 6th-grade students (ages 11-12) at a public middle school in the southeastern United States. The student population was 67% White, 14% Black/African American, 11% Asian, 8% Hispanic/Latino, and included three students of other races, with a near-equal gender distribution of 51% male and 49% female. The data collection and analysis protocol was approved by Vanderbilt University’s IRB, and all students whose data was analyzed provided consent. While 175 students participated in the study over two years, several students’ formative assessment responses were omitted from our analysis at various points due to either non-consent or absences resulting in incomplete work. In total, 158 instances were available for the Rules Task, 166 instances were available for the Debugging Task, and 161 instances were available for the Engineering Task.
@@
During Response Scoring, two of this paper’s authors independently scored a randomly selected subset (20%) of the data using the rubrics described in Section III. For each formative assessment, the two humans compared scores and discussed their disagreements before reaching a consensus. Cohen’s Kappa, the predominant measure in the literature assessing inter-rater reliability between two reviewers, was used as the IRR measure. The Cohen’s $\kappa$ values for the Rules, Debugging, and Engineering Tasks were 0.861, 0.740, 0.844, respectively.
@@
The datasets for all three tasks were imbalanced. In the Rules Task, the mode for total score was a perfect 9/9, occurring in 34 out of 158 cases, while a score of 0 was the second most frequent (observed in 29 instances). While examining the six subscores related to conditional statements and runoff values, a majority of students earned points. Conversely, for the three subscores related to absorption values, students were more likely not to receive points, suggesting students generally understood the logic expressions and runoff values but struggled to understand their relationship to absorption. The Debugging Task’s distribution was biased towards higher scores, with the mode being a perfect 5/5 for 66 out of 166 students and demonstrating a decreasing frequency of students attaining lower scores. All five Debugging Task subscores resulted in students earning points more often than not. The Engineering Task was predominantly characterized by incorrect responses, with the most common score being 0 (accounting for 63%; 101 out of 161). The remaining scores were evenly distributed across the range from 1 to 4 (inclusive).
@@
To evaluate CoTAL’s scoring accuracy on the Rules, Debugging, and Engineering Tasks, we compared the performance of CoTAL to a zero-shot, “scoring-only” Baseline (i.e., numerical scores only without labeled instances, CoT reasoning chains, or active learning) to evaluate CoTAL’s generalizability, comparing CoTAL-generated scores and explanations to those of the non-prompt-engineered LLM. Performance details for adding each individual component to CoTAL’s prompt engineering pipeline are reported in previous work [12]. We used GPT-4333A temperature of 0 was used to achieve near-deterministic behavior. to conduct our analysis due to its balance between cost and accuracy.
@@
To assess CoTAL’s generalizability for scoring and explaining formative assessment questions across multiple connected domains (RQ1), we adopted a mixed-methods approach. First, we measured LLM performance on a test set quantitatively, using Cohen’s QWK due to its prevalence in the automated essay scoring literature [39]. Unlike traditional Cohen’s $\kappa$, Cohen’s QWK awards partial credit based on the degree of disagreement, making it ideal for ordinal data. Next, we qualitatively investigated CoTAL’s impact on GPT-4’s scoring accuracy and its ability to provide useful feedback by conducting a constant comparative analysis of the LLM’s scoring explanations and errors to determine CoTAL’s strengths and weaknesses. CoTAL details for each of the three formative assessments are provided in our Supplementary Materials [^4] https://github.com/claytoncohn/TLT25_Supplementary_Materials, and we discuss our findings for RQ1 in Section V.
@@
To examine student and teacher impressions of CoTAL’s feedback efficacy and impact in supporting classroom learning (RQ2), we employed qualitative analysis by first conducting two semi-structured interviews with two classroom teachers. Each teacher had over five years of experience teaching the SPICE curriculum and more than 20 years of overall teaching experience. Teachers reviewed LLM responses to previously unseen Science Concepts and Reasoning (F1 in Figure 1) assessment questions, expressing their agreement with the LLM’s scores and explanations and sharing their preferences regarding the response structure. Additionally, we asked teachers how we could improve LLM outputs to further benefit both students and teachers.
@@
Second, we ran a separate focus group study with a subset of our study participants (23 students) to evaluate CoTAL’s performance in scoring their Science Concepts and Reasoning assessment. The students reviewed the AI-generated scores and feedback and then completed a survey. The survey assessed their agreement with GPT-4’s scoring accuracy, its feedback utility, and their confidence in the system’s ability to grade future assignments. To answer RQ2, we (1) created memos of key findings from the teachers’ interviews and (2) conducted a constant comparative analysis of the students’ survey responses. RQ2 results are presented in Section VI.
V Analyzing RQ1: CoTAL Generalizability Across Multiple Connected Domains

### v3

@@ 4.1 Chain-of-Thought Prompting + Active Learning (CoTAL)
CoTAL comprises three phases, as illustrated in Figure 3. Phase I (yellow) is human-driven: teachers and researchers apply ECD to develop curriculum-aligned formative assessments and rubrics, guided by the evidence needed to identify student mastery relative to curricular goals, and enabling effective evaluation of student responses (Section 3). The LLM is not involved in Phase I. Phase II (blue) co-develops and optimizes an initial prompt with the LLM by:
  1. 1.
    Response Scoring, which samples student responses, conducts IRR (e.g., Cohen’s Kappa [52]), and resolves disagreements to establish consensus;
  2. 2.
    Prompt Development, which employs few-shot CoT reasoning aligned to human consensus and structures the prompt with targeted prompt patterns; and
  3. 3.
    Active Learning, which iteratively refines the prompt on a validation set to align generations with human scoring preferences.
These map to the three blue diamonds in Figure 3. Phase III deploys the refined prompt in classrooms to score responses and generate explanations for assigned scores. The researchers then use teacher and student critiques to inform targeted refinements to questions, rubrics, and prompts without degrading LLM scoring performance. This procedure extends prior work on the Science Concepts and Reasoning Task [19] to three integrated assessments in science, computing, and engineering (Section 3).
@@
During Response Scoring (Phase II, Step 1), experts independently score a random subset of student responses with Phase I rubrics, and establish their initial degree of agreement based on an IRR metric (e.g., Cohen’s $\kappa$ [52]). When discrepancies arise between the scores assigned by the two reviewers, the underlying reasons for these differences are analyzed and resolved to reach a final consensus. Difficult-to-resolve discrepancies, termed “sticking points,” receive special attention during Prompt Development (Phase II, Step 2) to refine the prompt structure and guide the LLM toward alignment with the human scorers. Sticking points often emerge from “edge cases” where the reviewers interpret the same rubric differently. For example, one Engineering Task sticking point was whether a general understanding of design constraints (i.e., the student did not explicitly mention any constraint by name) qualified for 2 points (both reviewers ultimately agreed this was acceptable). This process—randomly sampling the data, calculating Cohen’s $\kappa$, and documenting reasons for disagreement—is repeated until $\kappa\geq 0.70$ to ensure consensus. This is depicted as the “IRR” self-loop in Figure 3.
@@
During Prompt Development (Phase II, Step 2), the LLM is first given task instructions, followed by the formative assessment question, grading rubric, and additional context to guide the LLM’s scoring. This supplementary information helps the LLM connect student responses to the rubric and comprehend the intended concepts, questions, and scoring criteria. Additionally, various prompt patterns [53] are employed to guide the LLM during inference and help structure its output, including a: (1) persona pattern that assigns the LLM a “persona,” or role, to play while generating its output; (2) context manager pattern that defines the context the LLM should consider while generating its output; (3) template pattern that provides the LLM with a structured output template; and (4) meta language creation pattern that creates a custom language for the LLM to understand (e.g., a textual shorthand notation to help the LLM interpret graphs) [53]. For example, the context manager pattern helped the LLM understand the relationship between the Rules and Debugging Tasks (i.e., the students often referred to conditional statements by “rule” number during the Debugging Task), and the meta language creation pattern allowed us to distill the Debugging Task’s computational model image into LLM-readable text.
@@
After constructing the prompt, few-shot instances are selected, equipped with CoT reasoning chains, and appended to the prompt to align the LLM with the consensus reached by human scorers. Each formative assessment prompt features two types of few-shot examples: (1) ground truth instances, where both scorers agreed during Response Scoring; and (2) sticking point instances, where scorers disagreed during Response Scoring, and the reasons for disagreement carried over to other instances, leading to similar disagreements among the scorers. For sticking point instances, CoT reasoning chains help clarify potential misunderstandings and guide the LLM toward human consensus. All few-shot instances are accompanied by CoT reasoning chains that loosely (i.e., not verbatim) follow this template:
@@
For instance, consider the following assessment question, rubric item, student response, and chain-of-thought from our Engineering Task (FA5) prompt: [^4] This example is used as a few-shot exemplar in the Engineering Task prompt. Prompts and demonstration instances for all three formative assessments are provided in our Supplementary Materials: https://github.com/claytoncohn/ijAIED25_Supplementary_Materials.
Question: A fictitious student, Morgan, has created two different designs. Morgan wants to test both of her designs to see which one is better [both are presented graphically]. Can Morgan tell which design is better based on these tests? Explain why or why not.
Rubric Item: 4 points: The student indicates the two designs cannot be compared because Morgan used different rainfall values to test each design AND the student explains that the tests are not fair (because runoff cannot be equally compared for the two tests if different amounts of rainfall are used). The student does not have to use the word “fair” explicitly to receive credit. He or she can indicate that the tests are not fair by mentioning that the two tests are uneven, inconsistent, impossible to compare, etc.
Student Response: No. Morgan has different inches of rainfall, which means that it is not equal or fair.
Chain-of-Thought: The student indicates he or she understands the two designs cannot be compared by providing an Answer of ‘No’. Additionally, the student says, ‘Morgan has different inches of rainfall’, which demonstrates the student understands Morgan used different rainfall values to test each design. The student also mentions that ‘it is not equal or fair’, which indicates the student understands that the tests are not fair. The student’s response, therefore, meets the Rubric criteria for a maximum score of 4 points. Based on the Rubric, the student earned a score of 4.
@@
Once the prompt is constructed, Active Learning (Phase II, Step 3) tests the prompt against a validation set. The LLM’s generations are analyzed by first isolating the incorrectly scored instances. For each subscore that the LLM predicts incorrectly, we identify a “scoring trend” to determine whether the LLM tends to produce false positives (FPs) or false negatives (FNs) and qualitatively discern the reasons behind these inaccuracies. This process mirrors Response Scoring, where recurring LLM errors on the validation set are identified as sticking points. We then select validation set instances exemplifying these sticking points, annotating them with CoT reasoning chains to correct the LLM’s mistakes and adding them to the existing few-shot examples in the prompt (illustrated by the “AL Prompt Refinements” loop in Figure 3).
@@
Phase II of CoTAL shares similarities with explainable AI (XAI) approaches, which focus on explaining a model’s decisions based on its internal mechanisms. However, in CoTAL’s case, the explainability comes from the LLM’s generations, offering insights into how the model evaluates student responses in relation to rubrics. The emphasis is on alignment with grading criteria rather than interpreting the model’s internal logic. This explainability is particularly valued by students, who see it as a critical factor for building trust in AI systems in educational settings (see Section 6). Unlike our prior work, which primarily highlighted our human-in-the-loop prompt engineering contributions, this paper “closes the loop” (see Figure 3) by (1) integrating LLM prompt engineering with ECD principles, (2) investigating our method’s practical utility in classrooms through studies involving both teachers and students, and (3) leveraging student and teacher feedback to inform refinements to our formative assessments, rubrics, and prompts.
@@
In this work, we refine our original Phase II approach based on prior findings [19]. Previously, we conducted active learning by inserting several validation instances back into the prompt. However, we found that the LLM tended to overfit when the number of few-shot instances was large, or the CoT reasoning chains became too granular. In this study, to mitigate the LLM’s tendency to overfit, we insert only a single instance into the prompt during Active Learning, specifically targeting the most persistent LLM errors. Additionally, we use Prompt Development (Phase II, Step 2) to provide the LLM with a list of “guidelines” [^5] In the actual prompt, we refer to the guidelines as “rules,” but we use the term “guidelines” in the manuscript so as not to be confused with the “rules” in the Rules Task. that the LLM is instructed to adhere to at all times. For example, one Rules Task guideline instructed the LLM not to consider the order in which students listed the three rules in its scoring decisions. Human scorers developed these guidelines based on the consensus they reached during the Response Scoring process. Just as we use CoT reasoning to cite evidence from a student’s response and connect it to the rubric for scoring, we also apply CoT reasoning to reference these guidelines in order to enhance the responses generated by the LLMs.
@@
In Phase III, the refined prompt from Phase II is implemented in a classroom setting to evaluate students’ responses to formative assessments and provide explanations for the assigned scores. Researchers then sample the responses generated by the LLM and present them to teachers for critique. The teachers assess the model’s scoring accuracy, soundness of its explanations, and clarity of its responses. During these discussions, the strengths and weaknesses of the language model are identified. Based on the feedback received, the teachers and researchers collaborate to address the model’s shortcomings while maintaining its strengths. They agree on specific refinements for the formative assessment questions, scoring rubrics, and prompts.
@@
Similarly, students are shown the LLM’s responses to their formative assessment answers and are asked to critique them. While teacher feedback informs methodological improvements related to curricular goals, student feedback highlights user experience and personalized learning. For example, students often emphasize elements such as the LLM’s tone in its responses and the overall effectiveness of the content it provides to enhance their understanding of relevant concepts. This process is illustrated in Figure 3 by the “Curricular and Instructional Refinements” and “Teacher/Student Prompt Refinements” loops. Like Phase I, Phase III depends on human input, involving close collaboration among researchers, teachers, and students to guide curricular and methodological enhancements.
In this study, we conducted Phase I (formative assessment and rubric development) as a participatory design effort between teachers and researchers. Phase II (prompt engineering) was carried out by members of our research team. Phase III involved continued participatory design with researchers, teachers, and students. This study presents findings on CoTAL’s scoring accuracy and its ability to generate scoring explanations (RQ1; Phase II), as well as insights from teachers and students to inform instructional and methodological refinements (RQ2; Phase III). In future work, we will examine the extent to which our Phase III refinements improve the system’s ability to meaningfully score student formative assessment responses and provide effective feedback to both students and teachers. The experimental procedure for this study is detailed in the following subsection.
4.2 Experimental Design
@@
We analyzed formative assessment responses from 175 sixth-grade students (ages 11-12) at a public middle school in the southeastern United States. The student population was 67% White, 14% Black/African American, 11% Asian, and 8% Hispanic/Latino, with three students identifying as other races. The gender distribution was 51% male and 49% female. The study protocol was approved by Vanderbilt University’s IRB, and all analyzed data were obtained from students who provided informed assent to participate and whose parents provided informed consent. Data were collected over two academic years. Although 175 students participated, some responses were excluded due to non-consent or absences, which resulted in incomplete work. Consequently, 158 students were available for the Rules Task, 166 for the Debugging Task, and 161 for the Engineering Task. The analytic sample, therefore, comprises all consenting, present sixth-grade students from the study site across a two-year period.
@@
During Response Scoring, two of this paper’s authors independently scored a randomly selected subset (20%) of the data using the rubrics described in Section 3. For each formative assessment, the two humans compared scores and discussed their disagreements before reaching a consensus. Cohen’s $\kappa$, the predominant measure in the literature assessing inter-rater reliability between two reviewers, was used as the IRR measure. The Cohen’s $\kappa$ values for the Rules, Debugging, and Engineering Tasks were 0.86, 0.74, 0.84, respectively.
@@
All three task datasets were imbalanced. In the Rules Task, the modal total score was a perfect 9/9, occurring in 34 out of 158 cases; 0 was the second most frequent score observed in 29 instances. Across the six subscores on conditional statements and runoff values, most students earned points; in contrast, for the three absorption-related subscores, students were more likely to receive no credit, indicating a stronger understanding of logic expressions and runoff values than their relationship to absorption. The Debugging Task’s distribution was biased towards higher scores, with the mode being a perfect 5/5 for 66 out of 166 students and demonstrating a monotonically decreasing frequency of students attaining lower scores. All five subscores showed students earning points more often than not. The Engineering Task was dominated by incorrect responses, with 0 as the most common score (63%; 101/161 students); the remaining scores were approximately uniformly distributed from 1 to 4 (inclusive).
@@
To evaluate CoTAL’s scoring accuracy on the Rules, Debugging, and Engineering Tasks, we compared the performance of CoTAL to a zero-shot, “scoring-only” Baseline (i.e., numerical scores only without labeled instances, CoT reasoning chains, or active learning) to evaluate CoTAL’s generalizability, comparing CoTAL-generated scores and explanations to those of the non-prompt-engineered LLM. Performance details for adding each individual component to CoTAL’s prompt engineering pipeline are reported in previous work [19]. We used GPT-4666A temperature of 0 was used to achieve near-deterministic behavior. to conduct our analysis due to its balance between performance and cost.
@@
To evaluate CoTAL’s generalizability for scoring and explaining formative assessment questions across connected domains (RQ1), we employed a mixed-methods design. Quantitatively, we assessed LLM performance on a held-out test set using Cohen’s QWK, selected for its prevalence in automated essay scoring [55]. Unlike Cohen’s $\kappa$, Cohen’s QWK awards partial credit proportional to the degree of disagreement, making it well-suited to ordinal labels and better aligned with educational rubrics (i.e., it mirrors how human graders value score proximity and partial correctness). Qualitatively, we conducted a constant comparative analysis of GPT-4’s scoring explanations and errors to characterize CoTAL’s impact on accuracy and feedback utility, and to identify strengths and weaknesses. Implementation details for each formative assessment are provided in Appendix A; prompts are available in the Supplementary Materials. We discuss our findings for RQ1 in Section 5.
@@
To investigate student and teacher perceptions of CoTAL’s feedback efficacy and its impact on classroom learning (RQ2), we employed qualitative analysis, analyzing teacher interviews and a student a sample of participating students. First, we conducted two semi-structured interviews with two classroom teachers, each with over five years of SPICE instructional experience and more than 20 years of overall teaching experience. Teachers reviewed LLM responses to previously unseen Science Concepts and Reasoning (F1 in Figure 1) assessment questions, stated their agreement with the LLM’s scores and explanations, and articulated preferences regarding response structure. They also provided recommendations for improving LLM outputs to support students and teachers better.
@@
For students, we conducted a focus group study with 23 consenting students, all present in one classroom during the second year of our two-year study, to evaluate CoTAL’s scoring of their Science Concepts and Reasoning assessment. The students reviewed the AI-generated scores and feedback and then completed a survey. The survey assessed their agreement with GPT-4’s scoring accuracy, its feedback utility, and their confidence in the system’s ability to grade future assignments. To answer RQ2, we (1) created memos of key findings from the teachers’ interviews and (2) conducted a constant comparative analysis of the students’ survey responses. RQ2 results are presented in Section 6.
5 Analyzing RQ1: CoTAL Generalizability Across Multiple Connected Domains